# Identifying Halos in Cosmological Simulations with Continuous Wavelet Analysis: The 2D Case

### v2

@@ 4.1 Determining Significance Levels and Threshold Criteria
In the previous section, we applied thresholding to the local maxima due to the presence of Poisson noise. The CWT of a uniform density field should ideally be zero, but the representation of the true density field using finite particles constitutes a Poisson process. This process leads to fluctuations in the CWT values, resulting in the generation of maxima and minima. Principally, there are two distinct methods to calculate the threshold when identifying structures using wavelets: one is based on the standard deviation of the data (e.g. Grebenev et al., 1995; Lazzati et al., 1999; Cayón et al., 2000; Barnard et al., 2004, and others), and the other employs Monte Carlo simulation (e.g. Bijaoui et al., 1992; Flin & Krywult, 2006; Pagliaro et al., 1999; Kazakevich et al., 2004). In this paper, we adopt the Monte Carlo method, as it allows us to pre-calculate the threshold for various cases. This approach yields an empirical formula that significantly simplifies the identification process.
The concept of employing random simulation to assess significance was initially introduced by (Bijaoui et al., 1992). This approach was subsequently adopted by (Slezak et al., 1993; Escalera et al., 1994) for the identification of structures using wavelets. The likelihood of identifying a spurious structure post-thresholding is termed the significance level. To evaluate the threshold, we can create a uniform Poisson noise ensemble that mirrors our data in particle count, as this noise originates from a Poisson process. It is crucial to recognize that each local maximum within the noise corresponds to a potential false structure. Consequently, following thresholding, any surviving grid points could result in a false identification. Therefore, the probability of a false identification, also referred to as the significance level, can be estimated by determining the proportion of CWT values that surpass the threshold.
To determine these thresholds, we conducted 800 iterations of Monte Carlo simulations at each scale, averaging the thresholds at a significance level of 0.005. Running Monte Carlo simulations each time we identify halos is time-consuming and computationally intensive. The threshold is dependent solely on the dimensionless scale parameter $k_{w}$ and the number of particles, assuming the resolution remains unchanged. Hence by performing Monte Carlo simulations for a specified resolution and across a wide range of values for both the particle number and $k_{w}$, we can derive an empirical formula. This formula will facilitate future halo identification processes.
We find that for any given particle number $N_{\rm p}$, out threshold grows linearly with $k_{w}$ (see left panel of the Figure 3). Thus, we assume our empirical formula has the following form:
@@
where $k$, $b$ represent the slope and intercept of the linear function, respectively, and both are solely dependent on $N_{\rm p}$. From the right panel of Figure 3, the slope and intercept can be fitted with a cubic function. Consequently, our empirical formula takes the following form:
@@
$k(N_{\rm p})=$ $3.0449\times 10^{-18}N_{\rm p}^{3}-1.3225\times 10^{-11}N_{\rm p}^{2}$
$+$ $1.9170\times 10^{-5}N_{\rm p}-0.8680,$
$b(N_{\rm p})=$ $6.2348\times 10^{-18}N_{\rm p}^{3}-3.1443\times 10^{-11}N_{\rm p}^{2}$
$+$ $4.3582\displaystyle 7.0298\times 10^{-5}N_{\rm p}-$ 4.3582 .
Equation (14) is valid within the range $50.1<k_{w}<$ 5 and $10^{5}<N_{\rm p}<3\times 10^{6}$, exhibiting an approximate dispersion of 10%. As $k_{w}$ increases, the support of each particle seldom intersects, causing the threshold to decrease with larger $k_{w}$ and to deviate from the linear relationship in scenarios with fewer particles.
Figure 9: The three largest FOF halos from the slices ${\rm z}=11.4,28.4$, and $38\,{h^{-1}{\rm Mpc}}$ are shown, along with their CWT contours. In dense regions, the CWT halo excludes the outermost parts of the FOF halo.
@@ 4.2 Optimal Grid Resolution for Accurate Halo Representation
To accurately represent the original CWT with a finite grid, it is preferable to sample the CWT on a grid that is not too sparse, given its approximate compact support. Each grid point can only sample particles within a finite area centered on the grid point. However, increasing the density of the grid points significantly increases computational costs. Therefore, we tested three sets of different resolutions:
@@
for intermediate resolution, and finally, for low resolution, we set
@@
We depict the same largest halo in a $38\,{h^{-1}{\rm Mpc}}$ slice across different resolutions to illustrate the resolution effect in Figure 4. We identified 1,059, 1,673, and 1,257 halos containing 164,885, 180,834, and 183,772 particles, respectively, from low to high resolution. For the largest halo, the low and high resolution images contain more particles (77,695 and 80,615) than the medium resolution (49,599), resulting in a lower count of halos detected in their runs. The low-resolution and high-resolution images exhibit a more elongated shape, whereas the medium-resolution image appears more compact. All three figures represent the same halo, and without a doubt, the high-resolution image provides the smoothest boundary.
@@
As defined in Section 3.3, the true structure corresponds to the largest CWT maxima across scales. So in both high-resolution and low-resolution runs, the largest halo is incorrectly selected due to the shift of the local maxima. Our cross-scale comparison is based on the positions of these maxima. If a maximum falls outside of our comparison range, it will still be considered valid, even if it does not accurately represent the true scale of the structure. The CWT maxima value of structures shown in the left panel and right panel of Figure 4 only reaches approximately $200\,{h^{-1}{\rm Mpc}}$, while the value of the true cross-scale CWT maxima is larger than $400\,{h^{-1}{\rm Mpc}}$.
Increasing the grid resolution typically separates closely linked structures that are obscured by low spatial sampling. However, in this particular case, doubling the resolution leads to the inclusion of even more particles in the largest halo, which is comparable to a FOF cluster containing 86,331 particles. What’s more, the shift of maxima causes the selection of larger structure as halo, which leads to a decrease in halo number. To prevent the halo from becoming excessively large and to reduce computational costs, we ultimately chose the intermediate resolution. At this resolution, we sample each individual CWT - specifically, just the MH wavelet itself - using more than 400 grid points within its support. This approach allows for an accurate representation of the CWT’s shape. Additionally, as can be seen from Figure 9, with a sufficiently high grid resolution, the intermediate image’s boundary most closely matches the true boundary.
Figure 10: Four individual CWT halos from the slice of ${\rm z}=38.0\,{h^{-1}{\rm Mpc}}$, and the linking length in this slice is $0.018\,{h^{-1}{\rm Mpc}}$. We can easily recognize that these particles would not be linked as a whole halo using the FOF method.
Table 1: The properties of FOF halo and CWT halo in our four slices.
@@ 5 Results
As previously mentioned, the clustering information is well-preserved in the CWT of the density field, as shown in Figure 5. This figure presents four different CWTs for the slice at ${\rm z}=38.0\,{h^{-1}{\rm Mpc}}$, covering scales from large to small. Notably, all four CWTs exhibit a spatial distribution similar to that of the CIC density. Higher densities in the density field also result in higher CWT values, while the absence of particles corresponds to a CWT value approaching zero. Additionally, every positive clump in the CWT is encircled by a high-negative ring caused by the matching of the negative ring of the wavelet and the density peak, which is the boundary of these clumps in our definition.
The figures confirm the ability of the CWT to distinguish structures across scales. The CWT at a larger scale contains larger clumps, and as the scale parameter increases (towards a smaller physical scale), the size of these clumps decreases. The CWT at any given scale reflects the clustering of particles into clumps at a size corresponding to this scale, and the value of the CWT represents the amplitude or strength of its corresponding clump.
@@
The detailed comparison between the FOF halo and the CWT halo are summarized in Table 1. The FOF method accounts for approximately 70% or less of the total particles, while the CWT accounts for over 60% of the particles. Specifically, the CWT method identifies halos containing 10% fewer particles in total and has a slightly higher number of halos compared to the FOF method. Moreover, the largest FOF halo contains approximately 30% more particles on average than that of CWT halo.
@@
When focusing on the densest regions, it is evident that the largest CWT halos are consistently surrounded by several intermediate-mass halos containing a few hundred particles. In contrast, the neighborhood of the largest FOF halo contains fewer particles, even in the final slice where the largest CWT and FOF halos are nearly identical. The discrepancy is primarily attributable to the compact shape of the CWT halo, although it is likely that the ‘linking bridge’ problem inherent in the FOF method also contributes. In dense regions, the high particle density promotes the formation of a chain of particles that links two nearby halos, resulting in a larger halo.
@@
Figure 8 presents the largest halos identified by the FOF method and their corresponding CWT halos. Each point represents a particle within this halo, and its color corresponds to the local density. These halos share similar positions and shapes, suggesting that they can be considered as the same halo classified by different methods. It should be noted that in the first three slices, the CWT halos are also identified as the largest halos. In the final slice, the largest FOF halo corresponds to the second largest CWT halo because it contains a similar number of particles.
@@
Figure 11: Halo mass function across four slices: the solid line represents the mass function of the CWT halos, and the dashed line represents that of the FOF halos. We can observe that the mass function of the FOF halos exhibits a bulge around a mass of 10 particles.
@@
In the comparison between CWT and FOF halos, it is observed that CWT halos have fewer particles, are more compact, and are defined by a clearer boundary at grid points where the CWT values are less than zero and decreasing from their local maximum. On the other hand, FOF halos are more diffuse, bounded by voids devoid of particles, and contain more particles. Both FOF and CWT halos feature several density peaks (represented by red particles) and display a multimodal structure indicative of the subhalos within them.
@@
The local density of CWT halos spans approximately two orders of magnitude, slightly less than that of FOF halos, whose local density spans $\lesssim 3$ orders of magnitude. This discrepancy is due to the no-particle area at the boundaries of FOF halos, which extend at least the linking length. In contrast, the boundary of CWT halos is defined by the change of the density field. As previously mentioned, the CWT of a constant background is zero. Therefore, our CWT halos are delineated solely based on the variation of the density field, resulting in sharp boundaries. This enables us to distinguish between different structures even in highly dense regions, identify each local density peak as a distinct halo, and effectively mitigate the linking bridge problem.
@@
Figure 8 shows that all four CWT halos exclude the outermost layer of their corresponding FOF halos and exhibit a slightly more compact density distribution. The compact shape of the CWT halo is attributed to the boundary condition we use. CWT equals zero means the the convolution of the negative part of wavelet with the central density peak is comparable to the convolution of the positive part of wavelet with outer part of the density peak. The high density peak forces the boundary of CWT equals zero inward to alleviate the the large negative convolution. Consequently, the diffuse and extended outermost parts are not included in the CWT halos. Figure 9 displays the largest FOF halo and its corresponding CWT contours. The scale of the CWT for these contours is set to $w\sim 4\,{h{\rm Mpc}^{-1}}$ for the first two halos and $w\sim 3\,{h{\rm Mpc}^{-1}}$ for the last halo, consistent with the scales of their respective CWT halos.
@@
The contours represent the boundary of the CWT halo with a high enough resolution that cannot be achieved by running the global halo identification program. The CWT halo exhibits a smooth boundary, contrasting with the fractal structure observed in FOF when the resolution is sufficiently high. Additionally, we affirm the importance of introducing our new boundary condition that delineates adjacent structures at the saddle point between them, as depicted in the final image of Figure 9. The large support of the wavelet at a large scale could create a positive bridge linking two adjacent halos if we were to use ‘CWT equals zero’ as the sole boundary condition.
@@
In regions of lower density, the finite resolution no longer poses a problem due to the sharp boundary, as the grid points that constitute the halo already encompass the no-particle voids. This implies that the uniform background density is close to zero in low-density regions. Figure 10 illustrates examples of these CWT halos. Within these halos, the distance between some groups of particles exceeds the linking length, leading to their corresponding FOF halo having fewer particles or even not forming at all due to an insufficient number of particles. This could partially account for why there are more CWT halos than FOF halos in the medium mass range.
@@
Additionally, the CWT method separates structures based on their CWT values, enabling the identification of relatively weaker and more extended structures within loose regions. These structures, while having a low CWT value, still exceed the threshold and are characterized by a negative ring surrounding them. Consequently, our CWT method can identify these particles as distinct individual halos.
@@
The density distribution of halos can be further examined through the halo mass function depicted in Figure 11. The function indicates that the number of CWT halos exceeds that of FOF halos in the very low mass range (less than 10 particles, near the minimum particle threshold), where CWT has an excess of over 100% in the number of halos. However, as the mass increases to approximately 10 particles, the number of CWT halos abruptly drops to about 40% of FOF. In the higher mass range, the number of CWT halos gradually approaches that of FOF halos and surpasses FOF when the mass reaches around 30 particles. Beyond this point, this excess stabilizes at approximately 20%, with a dispersion of about 20%. In the highest mass range (exceeding 2,000 particles), the scarcity of halo numbers introduces numerous artifacts and inaccuracies. There are even instances where no halos are present across many mass ranges, leading to discontinuities in our halo mass function when plotted on a logarithmic scale.
@@
Overall, all halo mass functions decrease as mass increases, but the CWT function has a smoother shape with no bulge at approximately 10 particles. This function demonstrates high consistency with the scatter observed among those halos where CWT indicates a higher abundance in the intermediate and lower mass ranges.
@@
Finally, we compute the halo power spectrum for four slices by assigning the halos to a $256^{2}$ grid using a weighted CIC method. The resulting power spectra are displayed in Figure 12. It is observed that the FOF halo exhibits higher amplitude across all wavenumbers due to containing more particles than the CWT halo. There is a high degree of consistency in the power spectrum across all scales. After a simple translation, the power spectra of the CWT and FOF halos nearly coincide. In summary, the clustering intensity of the CWT halo is similar to, but slightly less than, that of the FOF halo at every scale.

### v3

@@ 4.1 Determining Significance Levels and Threshold Criteria
In the previous section, we applied thresholding to the local maxima due to the presence of Poisson noise. While the CWT of a uniform density field should ideally be zero, the representation of the true density field using finite particles constitutes a Poisson process. This process induces fluctuations in the CWT values, resulting in the generation of spurious maxima and minima. Principally, there are two distinct methods to calculate the threshold when identifying structures using wavelets: one is based on the standard deviation of the data (e.g., Grebenev et al., 1995; Lazzati et al., 1999; Cayón et al., 2000; Barnard et al., 2004, and others), and the other employs Monte Carlo simulations (e.g., Bijaoui et al., 1992; Flin & Krywult, 2006; Pagliaro et al., 1999; Kazakevich et al., 2004). In this paper, we adopt the Monte Carlo method, which provides an empirical formula that significantly simplifies the identification process.
The concept of employing a random simulation to assess significance was initially introduced by Bijaoui et al. (1992). This approach was subsequently adopted by Slezak et al. (1993) and Escalera et al. (1994) for the identification of structures using wavelets. The likelihood of identifying a spurious structure post-thresholding is termed the significance level. To evaluate the threshold, we can create a uniform Poisson noise ensemble that mirrors our data in particle count. It is crucial to recognize that each local maximum within the noise corresponds to a potential false structure. Consequently, following thresholding, any surviving grid points could result in a false identification. Therefore, the probability of a false identification, also referred to as the significance level, can be estimated by determining the proportion of CWT values that surpass the threshold.
Figure 5: The largest halo within the slice at ${\rm z}=38.0\,{h^{-1}{\rm Mpc}}$, presented from left to right at low ($250500\times k_{w}+$ 250), medium ($5001000\times k_{w}+$ 500), and high ($10002000\times k_{w}+$ 1000) resolutions. The color represents the probability density of particles.
To determine these thresholds, we conducted 800 iterations of Monte Carlo simulations at each scale, averaging the thresholds at a significance level of 0.005. Running Monte Carlo simulations each time we identify halos is time-consuming and computationally intensive. The threshold is dependent solely on the dimensionless scale parameter $k_{w}$ and the number of particles, assuming the resolution remains unchanged. Hence, by performing Monte Carlo simulations for a specified resolution and across a wide range of values for both the particle number and $k_{w}$, we can derive an empirical formula. This formula will facilitate future halo identification processes.
We find that the threshold grows linearly with $k_{w}$ for a fixed particle number $N_{\rm p}$, as illustrated in the left panel of Figure 4. Thus, we assume our empirical formula has the following form:
@@
where $k$, $b$ represent the slope and intercept of the linear function, respectively, and both are solely dependent on $N_{\rm p}$. From the right panel of Figure 4, the slope and intercept can be fitted using a logarithmic function and a quadratic function, respectively. Consequently, our empirical formula takes the following form:
@@
$k(N_{\rm p})=$ $8.9326\times{\rm log}_{10}N_{\rm p}-44.0766,$
$b(N_{\rm p})=$ $-5.4418\times 10^{-12}N_{\rm p}^{2}+4.5586\times 10^{-5}N_{\rm p}$
-- $4.3582$
For improved accuracy at the low-particle end, curve fitting excludes results with particle numbers above $3\times 10^{6}$. Equation (14) is valid within the range $50.1<k_{w}<$ 5 and $10^{5}<N_{\rm p}<3\times 10^{6}$, exhibiting an approximate dispersion of 10%. As $k_{w}$ increases, the support of each particle seldom intersects, causing the threshold to decrease with larger $k_{w}$ and to deviate from the linear relationship in scenarios with fewer particles.
@@ 4.2 Optimal Grid Resolution for Accurate Halo Representation
To accurately represent the original CWT with a finite grid, it is preferable to sample the CWT on a grid that is not too sparse, given its approximate compact support. Each grid point can only sample particles within a finite area centered on the grid point. However, increasing the density of the grid points significantly increases computational costs. Consequently, we evaluated three distinct resolution sets:
@@
for intermediate resolution; and finally, for low resolution, we set
@@
We depict the same largest halo in a $38\,{h^{-1}{\rm Mpc}}$ slice across different resolutions to illustrate the resolution effect in Figure 5, with color indicating the probability density of particles. The identification yielded 1059, 1673, and 1257 halos containing 164,885, 180,834, and 183,772 particles, with computation times of 287.1, 357.4, and 1246.2 s for low, medium, and high resolutions, respectively. For the largest halo, the low and high-resolution images contain more particles (77,695 and 80,615) than the medium resolution (49,599), resulting in a lower count of halos detected in their runs. The low-resolution and high-resolution images exhibit more elongated shapes, whereas the medium-resolution image appears more compact. All three figures represent the same halo, and without a doubt, the high-resolution image provides the smoothest boundary.
Figure 6: The CWT halos of our pseudo-2D data. The slices are arranged in the order of ${\rm z}=11.4$, $28.4$, $38.0$, and $45.7\,{h^{-1}{\rm Mpc}}$ from top left to bottom right.
@@
As defined in Section 3.3, the true structure corresponds to the largest CWT maxima across scales. So, in both high-resolution and low-resolution runs, the largest halo is incorrectly selected due to the shift of the local maxima. Our cross-scale comparison is based on the positions of these maxima. If a maximum falls outside of our comparison range, it will still be considered valid, even if it does not accurately represent the true scale of the structure. The CWT maxima value of structures shown in the left and right panels of Figure 5 only reaches approximately $200\,{h^{-1}{\rm Mpc}}$, while the value of the true cross-scale CWT maxima is larger than $400\,{h^{-1}{\rm Mpc}}$.
Figure 8: The largest halos in Table 2 identified by the CWT (top) and the FOF (bottom) methods, arranged in the order of ${\rm z}=11.4$, $28.4$, $38.0$, and $45.7\,{h^{-1}{\rm Mpc}}$ from left to right. All particles within these halos are scattered and colored based on the probability density of particles.
Figure 9: The three largest FOF halos from the slices ${\rm z}=11.4,28.4$, and $38\,{h^{-1}{\rm Mpc}}$ are shown, along with their CWT contours. The color represents the probability density of particles. In dense regions, the CWT halo excludes the outermost parts of the FOF halo.
Increasing the grid resolution typically separates closely linked structures that are obscured by low spatial sampling. However, in this particular case, doubling the resolution leads to the inclusion of even more particles in the largest halo, which is comparable to a FOF cluster containing 86,331 particles. Moreover, the shift of maxima results in the selection of larger structures as halos, leading to a reduction in the number of halos and the loss of certain halos. For example, the fifth-largest halo in the $z=11.4\,{h^{-1}{\rm Mpc}}$ slice, the third-largest halo in the $z=28.4\,{h^{-1}{\rm Mpc}}$ slice, and the second- and sixth-largest halos in the $z=38.0\,{h^{-1}{\rm Mpc}}$ slice are no longer identified. Given that the introduction of the CWT boundary condition is expected to mitigate the ‘linking bridges’ problem inherent in the FOF method, it is anticipated that the CWT catalog will include a greater number of low-to-medium mass halos (15-3000 particles). In this range, the FOF method identifies 400–500 halos, while our intermediate-resolution result identifies approximately 500 halos. In contrast to our objective, both high- and low-resolution results exhibit a reduction in the number of halos in comparison to that of the FOF method.
To enrich the halo catalogs and minimize computational costs, we ultimately selected an intermediate resolution. At this resolution, we sample each individual CWT - specifically, just the MH wavelet itself - using more than 400 grid points within its support. This approach allows for an accurate representation of the CWT’s shape.
Table 2: The properties of FOF halos and CWT halos in our four slices.
@@
Figure 10: Four individual CWT halos from the slice of ${\rm z}=38.0\,{h^{-1}{\rm Mpc}}$, and the color represents the probability density of particles. The linking length in this slice is $0.018\,{h^{-1}{\rm Mpc}}$. Consequently, these halos must contain more particles than their FOF counterparts, given that the distance between some groups of particles exceeds the linking length.
@@ 5 Results
The detailed comparisons between the FOF halos and the CWT halos are summarized in Table 2. The FOF method accounts for approximately 70% or less of the total particles, whereas the CWT accounts for over 60% of the particles. Specifically, the CWT method identifies halos containing 10% fewer particles in total and yields a slightly higher number of halos compared to the FOF method. Moreover, the largest FOF halo contains approximately 30% more particles on average than the largest CWT halo.
@@
When focusing on the densest regions, it is evident that the largest CWT halos are consistently surrounded by several intermediate-mass halos containing a few hundred particles. In contrast, the neighborhood of the largest FOF halo contains fewer particles, even in the final slice, where the largest CWT and FOF halos are nearly identical. This discrepancy is primarily due to the compact shape of the CWT halo, although it is likely that the ‘linking bridge’ problem inherent in the FOF method also contributes. In dense regions, the high particle density promotes the formation of a chain of particles that links two nearby halos, resulting in a larger halo.
@@
Figure 8 presents the largest halos identified by the FOF method and their corresponding CWT halos. Each point represents a particle within this halo, and the color represents the probability density of the particles. These halos share similar positions, shapes, and even particle content, suggesting that they can be considered as the same halo classified by different methods. It should be noted that in the first three slices, the CWT halos are also identified as the largest halos. In the final slice, the largest FOF halo corresponds to the second-largest CWT halo because it contains a similar number of particles.
@@
Figure 11: Top: the cumulative halo mass function across the four slices mentioned in Table 2, the solid line represents the cumulative mass function of the CWT halos, and the dashed line represents that of the FOF halos. Bottom: the relative difference between the cumulative halo mass function of FOF and CWT.
@@
In the comparison between CWT and FOF halos, it is observed that CWT halos have fewer particles, are more compact, and are defined by a clearer boundary at grid points where the CWT values are less than zero and decreasing from their local maximum. On the other hand, FOF halos are more diffuse, bounded by voids devoid of particles, and contain more particles. Both FOF and CWT halos feature several density peaks (represented by red particles), revealing a multimodal structure indicative of the subhalos within them.
@@
The local density of CWT halos spans approximately two orders of magnitude, slightly less than that of FOF halos, whose local density spans $\lesssim 3$ orders of magnitude. This discrepancy is due to the no-particle area at the boundaries of FOF halos, which extends at least the linking length. In contrast, the boundary of CWT halos is defined by the change in the density field. As previously mentioned, the CWT of a constant background is zero. Therefore, our CWT halos are delineated solely based on the variation of the density field, resulting in sharp boundaries. This enables us to distinguish between different structures even in highly dense regions, identify each local density peak as a distinct halo, and effectively mitigate the linking bridge problem.
@@
All four CWT halos exclude the outermost layer of their corresponding FOF halos, demonstrating a marginally more compact density distribution. The compact shape of the CWT halo is attributed to the boundary condition we use. CWT equals zero implies that the convolution of the negative part of the wavelet with the central density peak is comparable to the convolution of the positive part of the wavelet with the outer part of the density peak. The high-density peak forces the boundary where CWT equals zero inward to alleviate the large negative convolution. Consequently, the diffuse and extended outermost parts are excluded from the CWT halos. Figure 9 displays the largest FOF halos and their corresponding CWT contours. The scale of the CWT for these contours is set to $w\sim 4\,{h{\rm Mpc}^{-1}}$ for the first two halos and $w\sim 3\,{h{\rm Mpc}^{-1}}$ for the last halo, consistent with the scales of their respective CWT halos.
@@
The contours represent the boundary of the CWT halo with a high enough resolution that cannot be achieved when running the global halo identification program. The CWT halo exhibits a smooth boundary, contrasting with the fractal structure observed in FOF when the resolution is sufficiently high. Additionally, we affirm the importance of introducing our new boundary condition that delineates adjacent structures at the saddle point between them, as depicted in the final image of Figure 9. The extensive support of the wavelet at a large scale could create a positive bridge linking two adjacent halos if we were to use ‘CWT equals zero’ as the sole boundary condition. Moreover, it can be seen that with a sufficiently high grid resolution, the boundary of the intermediate-resolution image most closely matches the true boundary.
@@
In low-density regions, the finite resolution no longer poses a problem due to the sharp boundary, as the grid points that constitute the halo already encompass the no-particle voids. This implies that the uniform background density is close to zero in low-density regions. Figure 10 illustrates examples of these CWT halos. Within these halos, the distance between some groups of particles exceeds the linking length, leading to their corresponding FOF halo having fewer particles or even not forming at all due to an insufficient number of particles. This could partially account for why there are more CWT halos than FOF halos in the medium mass range.
@@
Furthermore, the CWT method separates structures based on their CWT values, enabling the identification of relatively weaker and more extended structures within loose regions. These structures, while having a low CWT value, still exceed the threshold and are characterized by a negative ring surrounding them. Consequently, our CWT method can identify these particles as distinct individual halos.
@@
The density distribution of halos can be further examined through the cumulative halo mass functions shown in Figure 11. These functions show that the number of CWT halos exceeds that of FOF halos in the very low mass range (less than 10 particles, near the minimum particle threshold), where CWT shows a relative excess of over 50%. However, as the mass increases to approximately 10 particles, the number of CWT halos drops sharply, creating a valley in the relative difference. In the higher mass range, the number of CWT halos gradually approaches that of FOF halos and surpasses FOF when the mass reaches approximately 30 particles, at which point the relative difference begins to grow. Beyond this threshold, this excess stabilizes at approximately 20%, with a dispersion of about 20%.
@@
Overall, the cumulative halo mass functions exhibit a decelerating growth rate as mass increases, indicating a declining trend in the mass function. The abrupt decline observed in the relative difference around 10 particles demonstrates a high degree of consistency with the scatter observed among those. Specifically, the CWT method indicates a higher abundance in the intermediate and lower mass ranges.
@@
Finally, we compute the halo power spectrum for four slices by assigning the halos to a $256^{2}$ grid using a weighted CIC method. The resulting power spectra are displayed in Figure 12. It is observed that the FOF halos exhibit higher amplitude across all wavenumbers due to containing more particles than the CWT halos. There is a high degree of consistency in the power spectrum across all scales. After a simple translation, the power spectra of the CWT and FOF halos nearly coincide. In summary, the clustering intensity of the CWT halos is similar to, but slightly less than, that of the FOF halos at every scale.